I have an algebra project and one of the questions says " The difference between two perfect squares is 133.What is the smallest possible sum of the two perfect squares? " I found out that the answer is 205, but I don't understand why that is the answer.
Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

I have an algebra project and one of the questions says " The difference between two perfect squares is 133.What is the smallest possible sum of the two perfect squares? " I found out that the answer is 205, but I don't understand why that is the answer.
------
Equation:
x^2 - y^2 = 133
(x-y)(x+y) = 133
-------
133 = 7*19 = (13-6)(13+6)
--------
Ans: 13^2+6^2 = 169 + 36 = 205
===================================
Cheers,
Stan H.
==============

Answer by Edwin McCravy(20054)

       x² - y² = 133

(x - y)(x + y) = 133

So (x - y) and (x + y) are factors of 133
and (x - y) is the smaller factor and (x + y) 
is the larger factor. 

The only two ways to factor 133 are

1·133 where 1 is the smaller and 133 is the larger.  

and

7·19  where 7 is the smaller and 19 is the larger.

In the first case we have

x - y = 1
x + y = 133

Solve that system and get solution x=67, y=66

The sum of their squares is 67²+66² = 4489 + 4356 = 8845 

--------------

In the second case we have

x - y = 7
x + y = 19

Solve that system and get solution  x=13, y=6

The sum of their squares is 13²+6² = 169 + 36 =  205

Edwin